OpenStudy (gabylovesyou):

Is this correct??? The volume of a cone is 300 cubic kilometers. If the radius of the base is 9 kilometers, what is the height of the cone? Use 3.14 for TT. Show your work and round to the nearest tenth. my answer ---> V = πr^2h 300 km^3 = 3.14(7 km)^2 h 1.949824516... = h The height is approximately 1.9 km. DID I DO IT RIGHT? lol

jimthompson5910 (jim_thompson5910):

You are very very close

jimthompson5910 (jim_thompson5910):

The volume of a cone is really V = (pi*r^2*h)/3

jimthompson5910 (jim_thompson5910):

So this means V = (pi*r^2*h)/3 300 = (pi*9^2*h)/3 300 = (3.14*81*h)/3 300 = (254.34h)/3 300*3 = 254.34h 900 = 254.34h 900/254.34 = h 3.53857 = h h = 3.53857 km

OpenStudy (gabylovesyou):

so the height would be approximately 3.5 km

jimthompson5910 (jim_thompson5910):

yes, if you're rounding to the nearest tenth

jimthompson5910 (jim_thompson5910):

which I just now saw as part of the instructions lol

OpenStudy (gabylovesyou):

ohhh ok!

jimthompson5910 (jim_thompson5910):

so yes, you have it

OpenStudy (gabylovesyou):

Thank you soo much!

jimthompson5910 (jim_thompson5910):

you're welcome